# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

### v2

@@ Abstract
Recently, discrete diffusion language models have demonstrated promising results in NLP. However, there has been limited research on integrating Pretrained Language Models (PLMs) into discrete diffusion models, resulting in underwhelming performance in downstream NLP generation tasks. This integration is particularly challenging because of the discrepancy between step-wise denoising strategy of diffusion models and single-step mask prediction approach of MLM-based PLMs. In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates PLMs with the diffusion models. Furthermore, as it is challenging for PLMs to determine where to apply denoising during the diffusion process, we integrate an entropy tracking module to assist them. Finally, we propose entropy-based noise scheduling in the forward process to improve the effectiveness of entropy-adaptive sampling throughout the generation phase. Experimental results show that Diffusion-EAGS outperforms existing diffusion baselines in downstream generation tasks, achieving high text quality and diversity with precise token-level control. We also show that our model is capable of adapting to bilingual and low-resource settings, which are common in real-world applications.
PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling
@@ 1 Introduction
Type Dataset OpenWebText Gokaslan and Cohen (2019) RocStories Mostafazadeh et al. (2016) Deontology Hendrycks et al. (2023) Question Generation Dhingra et al. (2017) QQP Wang et al. (2017) ALMA Xu et al. (2024a) ParaDetox Logacheva et al. (2022)
Open-ended Generation $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
Conditional Generation $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
– Context Provided ? - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
– Content Provided ? - $\times$ $\triangle$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
– Format Provided ? - - $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Table 1: Each dataset has a different level of conditional contraints even if they are conditional generation tasks. $\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP, and demonstrate their promising performance over autoregressive models (Li et al., 2022; Gong et al., 2023a; He et al., 2023; Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023; He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024; Nie et al., 2024). Moreover, these models have been shown to be scalable, and are able to address the challenge of generalizing bidirectional relationships while maintaining sensitivity to the temporal alignment between training and test data, a conventional architectural limitation often observed in autoregressive models (Nie et al., 2024).
Although recent Discrete Diffusion Language Models (DDLMs) have demonstrated high performance in NLP tasks such as unconditional and open-ended generation compared to Continuous Diffusion Language Models (CDLMs) (Lou et al., 2024), existing DDLMs continue to face limitations in conditional generation, thereby restricting their broader applicability across diverse NLP domains. Our experiments suggest that models such as SEDD falls short in such tasks.
To enhance performance of DDLMs, PLMs can be integrated into diffusion models. This approach leverages the capability of PLMs to elicit improved generalization capabilities in unseen tasks while simultaneously exploiting the controllability and diversity strengths inherent to diffusion models. However, integrating PLMs into diffusion models is non-trivial as PLMs typically predict masked elements in a single step, whereas diffusion models require step-wise denoising based on the overall semantics of each timestep sequence, and such gap yields limited results. Therefore, we need to consider such inconsistencies to effectively adopt PLMs into DDLMs by a new methodology.
In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for conditional generation. To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by adopting PLMs as the denoising function at each step of the diffusion process. Since identifying where to apply denoising during the diffusion process is challenging for PLMs in our experiments, we incorporate an entropy tracking module to support their operation based on entropy information. To train DDLMs to effectively exploit an entropy tracking module, we propose an entropy-based noising scheduling during training phase in a forward process. Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module.
Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to baseline models across various conditional generation tasks. Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, and validate that our model can also adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications, indicating the potential applicability of our model across a wide range of conditional generation tasks.
2 Task Setting
2.1 Fine-Grained Conditional Generation
In conditional generation tasks, the level of conditional constraint imposed by the dataset plays a critical role in shaping the generation process. As shown in Table 1, conditional constraints are diverse across datasets. In our task, we categorize these constraints into three levels: (1) the provision of context alone, requiring the continuity of the prefix; (2) the provision of specific content to be included in the target sequence, necessitating the inclusion of certain keywords; and (3) the provision of semantic content formatting, such as transforming toxic sentences into safer alternatives or converting text from the source language to a target language. In our study, we aim to develop a universal diffusion framework capable of being applied across a wide range of conditional generation tasks.
2.2 DDLM with BERT
BERT as Markov Random Field (MRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models. Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while PLMs are trained to predict masked elements in a single step. To effectively integrate PLMs into DDLMs, it is necessary to resolve such a discrepancy. Therefore, we begin by interpreting the intrinsic characteristics of MLMs as Markov Random Field (MRF), as demonstrated in Wang and Cho (2019). Details are in Appendix B.
constrained MRF (cMRF) Even though open-ended MLMs are MRFs, we observe a significant increase in log-potential values for sequences when guided by the RocStories conditions, as shown in Figure 1. Additional experiments supporting this observation are detailed in Appendix A.
This suggests that conditional generation models differ from open-ended models in randomness, making it crucial to investigate whether conditionally constrained spaces adhere to MRF properties. Through the derivation process outlined in Appendix B, we demonstrate that these spaces retain the characteristics of an MRF.
Given a sequence of fully connected random variables $X=(x_{1},x_{2},\ldots,x_{L})$ with a set of observed variables $Y=(y_{1},y_{2},\ldots,y_{K})$, $f_{\theta}(X\backslash\{x_{l}\};Y)$ can be represented by logits from MLM given the sequence $X\backslash\{x_{l}\}$ and the observed variables $Y$. Then, log-potential can be calculated as:
$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[MASK]}\notin X_{l-1:l+1},\\ \\ 0,&\text{otherwise}\end{cases}$
where $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1. By this, we can define the potential function for the clique as:
$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$ (1)
Then, the constrained probability of sequence X is:
$p_{\theta}(X;Y)=\frac{1}{Z(Y,\theta)}\prod_{l=1}^{L}\phi_{l}(X;Y)$
Using the softmax function, the distribution of a variable $x_{l}$ given the observed variables $Y$ is:
$p(x_{l}|X\backslash\{x_{l}\},Y)=\frac{\exp(1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y))}{\sum_{m=1}^{M}\exp(1h(m)^{T}f_{\theta}(X\backslash\{x_{l}\};Y))}$ (2)
This formulation implies that the dataset-constrained generation process in BERT can still be interpreted as a MRF with constraint Y, where the log-potential functions can be derived from the output logits of MLM.
3 Methodology
Figure 2: Given a conditional constraint $Y$ and a generation target $X$, the training process utilizes Entropy-based Sampling (ENS) to determine which positions in the masked sequence should be denoised when transitioning from one timestep to the next. The tokens at the corresponding masked positions ([M] denotes masked tokens) are then generated using $f_{\theta}$. The loss is subsequently computed, and the parameters are updated accordingly. During the inference process, the sequence $X$, initially composed of mask tokens, is iteratively refined through the diffusion process to generate $X$ that satisfies the given condition $Y$.
In this section, we propose Diffusion-EAGS that integrates PLM into DDLMs through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS). The overview of our process is in Figure 2.
3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
As discussed in Section 2.2, MLM-based PLMs can be interpreted as cMRFs, allowing Gibbs sampling to be applied at each denoising step in DDLMs by leveraging the properties of MRFs. As PLMs often struggle to select the next denoising targets, it is important to highlight that the primary purpose of adopting Gibbs sampling is not only to bridge the gap between PLMs and diffusion models but also to assist PLMs in identifying which absorbing tokens should be denoised. Therefore, we propose Entropy-Adaptive Gibbs Sampling (EAGS) to address the challenge PLMs face in identifying the appropriate denoising targets. In EAGS, the entropy of each variable quantifies the amount of uncertainty at a specific position within the sequence. This enables sampling of each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$ in the descending order of entropy. By prioritizing the generation of the least informative parts of the sequence, EAGS facilitates the creation of more structured sequences by leveraging the syntactic context that has already been established.
Specifically, according to Equation (2), the probability $p^{t}(x_{l}=v_{m}|Y,f_{\theta})$ of the token $v_{m}$ given the context $Y$ and $f_{\theta}$ at step $t$ can be calculated as:
$p^{t}(x_{l}=v_{m}|Y,f_{\theta})=\frac{\exp(f_{\theta}(v_{m}|X^{t}))}{\sum_{m^{\prime}=1}^{M}\exp(f_{\theta}(v_{m^{\prime}}|X^{t}))}$
, where $f(v_{m}|X^{t})$ is the logit corresponding to the token $v_{m}$ for the position $x_{l}$ at step $t$. That is, we can compute the conditional probability for each masked position. With the conditional probability, the entropy $H$ for a variable $x_{l}$, logits function $f_{\theta}$, and the dataset context $Y$ is derived as:
$\begin{split}&H(x_{l}|Y,f_{\theta})=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\log p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\end{split}$
With $H$, our approach for T step-generation process can be formalized as follows:
@@ 3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f_{\theta})$ for each variable $x_{l}$.
@@
    Variable Selection: Select the variable $x_{l^{*}}$ with the highest entropy for sampling
    $l^{*}=\arg\max_{l}H(x_{l}\mid Y,f_{\theta})$

### v3

# Conditional [MASK] Discrete Diffusion Language Model

@@ Abstract
Although auto-regressive models excel in natural language processing, they often struggle to generate diverse text and provide limited controllability. Non-auto-regressive methods could be an alternative but often produce degenerate outputs and exhibit shortcomings in conditional generation. To address these challenges, we propose Diffusion-EAGS, a novel framework that integrates conditional masked language models into diffusion language models through the theoretical lens of a conditional Markov Random Field. In doing so, we propose entropy-adaptive Gibbs sampling and entropy-based noise scheduling to counterbalance each model’s shortcomings. Experimental results show that Diffusion-EAGS outperforms baselines and achieves the best quality-diversity tradeoff, demonstrating its effectiveness in non-autoregressive text generation.
Conditional [MASK] Discrete Diffusion Language Model
@@ 1 Introduction
Figure 1: Overview of how our approach (Diffusion-EAGS) combines the strengths of MLM and diffusion-based models to overcome the limitations of AR models, achieving a better diversity-quality tradeoff and fine-grained controllability
Auto-Regressive Models (ARMs) have driven significant advances in NLP Achiam et al. (2023); Dubey et al. (2024); Team et al. (2023), yet they still face fundamental challenges such as diversity and controllability due to the ARM’s innate inductive bias. To address these challenges, a more flexible generative model is required.
Specifically, ARMs face multiple challenges: they struggle to correct mathematical reasoning errors once made Wang et al. (2025), and often fail to integrate external knowledge Hudecek and Dusek (2023); Sun et al. (2023); Su et al. (2024). These shortcomings arise from ARMs’ sequential nature, which prevents them from revising earlier steps. As a result, they cannot effectively foster diversity through temperature-based sampling alone Lee et al. (2025), nor can they anticipate future requirements at earlier steps, thus undermining controllability when specific keywords must appear later Lu et al. (2022).
One promising alternative is non-autoregressive generation, including Conditional Masked Language Models (CMLMs) Ghazvininejad et al. (2019a); Kasai et al. (2020) and diffusion models. CMLMs provide strong contextual understanding but lack an effective text generation mechanism. Meanwhile, diffusion models iteratively refine text through denoising, enabling fine-grained control and increased diversity. Recent works explore direct diffusion-based generation Li et al. (2022); Gat et al. (2024); The et al. (2024); Ye et al. (2025) or hybrid approaches combining diffusion with PLMs and LLMs (Lin et al., 2023; Xiang et al., 2024). However, despite their advantages, Discrete Diffusion Language Models (DDLMs) still suffer from degeneration in conditional generation tasks Xu et al. (2025), as confirmed by our experiments.
We therefore propose Diffusion-EAGS, a novel approach that integrates CMLMs into DDLMs to achieve diverse, controllable, and high-quality conditional generation. However merging these methods is challenging because CMLMs generate text in one step by predicting all masked tokens, whereas diffusion models iteratively refine representations over multiple steps by introducing and removing noise. Our approach bridges this gap by leveraging a conditional Markov Random Field (cMRF) formulation, which enables:
@@
    Stepwise iterative generation, overcoming the single-step limitations of CMLMs.
@@
    Stable and diverse conditional text generation, reducing semantic drift in DDLMs.
Diffusion-EAGS achieves this through two key methodologies:
  • Entropy-Adaptive Gibbs Sampling (EAGS): A strategy that updates the most uncertain (high-entropy) tokens first at each denoising step, ensuring qualified generation.
  • Entropy-based Noise Scheduling (ENS): A training approach that progressively masks tokens based on ascending order of entropy, enabling the model to learn a structured denoising process.
We conduct extensive experiments to validate Diffusion-EAGS on various conditional generation tasks, demonstrating significant improvements over baseline models. Our approach achieves the best quality-diversity tradeoff, demonstrating that Diffusion-EAGS balances fluency and variability more effectively than existing models. Moreover, keyword-based story generation experiments confirm that our model effectively generates coherent and controlled text from randomly masked sequences, making it highly adaptable to different conditioning constraints.
2 Related Works
Efforts to integrate generative flow models into sequence generation exploit the distribution shift from a source language to a target language through a series of invertible linear transformations (Ma et al., 2019; Zhang et al., 2024). However, as DDPM (Ho et al., 2020a) demonstrate the effectiveness of generating images, diffusion models have been a major topic of interest within the field of generative flow models (Song et al., 2021a, b). To apply such diffusion methodologies to NLP, in order to leverage their strengths in controllability and diversity, recent studies have demonstrated promising performance across various tasks (Li et al., 2022; Gong et al., 2023a; He et al., 2023; Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023; He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024; The et al., 2024; Wang et al., 2024).
Although Continuous Diffusion Language Models (CDLMs) such as Diffusion-LM (Li et al., 2022), DiffuSeq-v1, v2 (Gong et al., 2023a, b), and LD4LG (Lovelace et al., 2023) show promising performance, Bansal et al. (2022) argue that such operations do not necessarily have to be governed by stochastic randomness.
Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al. (2022) adopt pre-trained language models (PLMs) to DDLM. SEDD Lou et al. (2024) propose score entropy inspired by MLM loss, and outperform existing CDLMs. Recent works by Shi et al. (2024) and Sahoo et al. (2024) extend this idea and obtain better empirical results. Zheng et al. (2024) further enhance discrete diffusion models by correcting the numerical precision error in SEDD-based models. These research make an improvement on the open ended generation task. Furthermore, Venkatraman et al. (2024) use SEDD as text infilling, and Nie et al. (2024) demonstrate that DDLMs are scalable.
3 MLM & DDLM : D-cMRF
Pre-trained MLMs offer rich, context-aware representations through one-pass masked prediction, whereas DDLMs iteratively refine text via stepwise denoising to enhance control and diversity. Combining these approaches can overcome MLMs’ one-pass limitations and DDLMs’ degeneration in conditional generation. However, their integration is challenging because DDLMs require iterative updates while MLMs predict all masked tokens simultaneously. To bridge this gap, we propose Diffusion-based Constrained Markov Random Fields (D-cMRF), a framework that integrates a discrete diffusion process into MLM sequence generation. By leveraging an entropy-based sampling strategy to selectively update high-uncertainty tokens at each step, D-cMRF achieves a principled reduction in sequence energy, leading to stable and coherent generation.
3.1 MLM as cMRF
Inspired by the traditional approaches of Wang and Cho (2019) and Goyal et al. (2022), which model MLMs as Markov Random Fields (MRFs) and energy-based models (EBMs), respectively, we reinterpret MLM as a conditional MRF (cMRF) model and employ it as a denoising function at each diffusion step.
Let $X=(x_{1},x_{2},\dots,x_{L})$ be a sequence of discrete variables from a vocabulary $V$, with $Y$ representing observed conditions. The sequence probability follows an energy-based MRF formulation:
$P_{\theta}(X;Y)=\frac{\exp(-E_{\theta}(X;Y))}{Z(Y,\theta)}$ (1)
where $E_{\theta}(X;Y)$ is the energy function parameterized using MLM-based embeddings,  $\theta$ denotes parameterization of MLM, and $Z(Y,\theta)$ is the partition function for ensuring proper normalization. Then the total sequence energy is defined as:
$E_{\theta}(X;Y)=-\sum_{l=1}^{L}\log\phi_{l}(X;Y)$ (2)
where log-potential function $\log\phi_{l}(X;Y)$ is :
$\log\phi_{l}(X;Y)=1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y)$ (3)
where $l$ is a token position in the sequence, $1h(x_{l})$ is the one-hot encoding of token $x_{l}$, and $f_{\theta}(X\backslash\{x_{l}\};Y)$ represents the MLM logit output conditioned on the sequence.
3.2 DDLM with Entropy-based Denoising
Determining how to perform sampling with such a simple cMRF presents a separate challenge. In particular, one can use techniques such as Gibbs sampling as long as the energy space remains unchanged, but we cannot guarantee that this energy space is stable in general Goyal et al. (2022). The necessity of generating sequences in cMRF based on energy update is in Appendix A. Hence, a natural research question arises: “How should we sample and update the energy?”
The training process of diffusion models (both forward and backward) conceptually represents the entire distribution as a product of local conditional distributions across time steps. Hence, diffusion models share a probabilistic graphical structure with MRF, enabling MLM to be integrated within DDLM framework.
Therefore, in this subsection, we describe how to update the energy and perform sampling under the DDLM framework using $P_{\theta}(X;Y)$. Specifically, we integrate $P_{\theta}(X;Y)$ into each diffusion step as a denoising function, employing an entropy-based denoising matrix $Q$ in Section 4.2. We first define the entropy of each token:
$H_{i}(X^{(t)})=-\sum_{x^{\prime}\in V}p_{\theta}(x^{\prime}_{i};X^{(t)})\log p_{\theta}(x^{\prime}_{i};X^{(t)})$ (4)
where $p_{\theta}(x^{\prime}_{i};X^{(t)})$ is the softmax probability of token $x^{\prime}_{i}$ at position $i$ in sequence $X^{(t)}$, and $t$ denotes the diffusion timestep.
$M_{t}=\{i\mid H_{i}(X^{(t)})\geq\tau_{t}\}$ (5)
where $\tau_{t}$ is a dynamically adjusted entropy threshold. This ensures that updates occur at positions where the model has the highest uncertainty. Subsequently, we sample the next-step sequence from $P_{\theta}(X^{(t)};Y)$ at the suggested positions. We perform this selection process at every diffusion step, which corresponds to updating the energy, different from existing research Wang and Cho (2019); Goyal et al. (2022).
3.3 D-cMRF
By combining DDLM with cMRF, our approach enables a theoretically grounded generation process from the perspective of MLM. Moreover, from the diffusion standpoint, the training process naturally aligns with MLM objective, as discussed in Section 3.1 and Section 3.2. Specifically, our D-cMRF guarantees energy reduction during generation, ensuring stable sequence reconstruction.
Figure 2: Overview of the training (forward) and inference (backward) processes in Diffusion-EAGS. Training (left): Entropy-based Noise Scheduling (ENS) determines which tokens in the masked sequence, denoted by $[M]$, should be denoised at each timestep based on the position entropy $H(x_{i})$. These tokens are then generated using the diffusion model with parameters $\theta$, and the loss is computed using a cross-entropy (C.E.) diffusion loss. Inference (right): Starting from a fully masked sequence conditioned on $Y$, Entropy-Adaptive Gibbs Sampling (EAGS) iteratively refines the sequence by focusing on high-entropy tokens, denoted as $M_{t}$, based on a threshold $\tau_{t}$, yielding stable and coherent text generation.
Step 1: Expected Energy at Diffusion Step $t$
At diffusion step $t$, the expected sequence energy is defined as follows:
$\mathbb{E}_{X^{(t)}\sim q}\left[E_{\theta}(X^{(t)};Y)\right]=\sum_{X^{(t)}}q(X^{(t)})E_{\theta}(X^{(t)};Y)$ (6)
where $q(\cdot)$ denotes the probability distribution from which $X^{(t)}$ is sampled. Since high-entropy tokens are selected for replacement, the total sequence energy can be decomposed as follows:
$\mathbb{E}\bigl{[}E_{\theta}(X^{(t)};Y)\bigr{]}$ $=\sum_{i\in M_{t}}\mathbb{E}\bigl{[}E_{\theta}(x_{i}^{(t)};X^{(t-1)},Y)\bigr{]}$ (7)
$\quad+\sum_{i\notin M_{t}}E_{\theta}\bigl{(}x_{i}^{(t-1)};Y\bigr{)}.$
Step 2: Energy Reduction via Denoising
Since masked tokens are replaced with lower-energy candidates at each step, we expect a general trend of energy reduction. However, due to the stochastic nature of sampling, local fluctuations in energy may occur. Over multiple diffusion steps, the entropy-based selection mechanism ensures a net decrease in sequence energy.
$\mathbb{E}\left[E_{\theta}(x_{i}^{(t)};X^{(t-1)},Y)\right]\leq E_{\theta}(x_{i}^{(t)};X^{(t)},Y)$ (8)
Applying this property across all updated tokens $i\in M_{t}$, we obtain:
$E_{\theta}(X^{(t-1)};Y)\leq E_{\theta}(X^{(t)};Y)$ (9)
Step 3: Convergence to Low-Energy States
Summing over all diffusion steps $T$:
$E_{\theta}(X^{(0)};Y)\leq E_{\theta}(X^{(T)};Y)$ (10)
where $X^{(T)}$ is the fully masked sequence with maximum entropy, and $X^{(0)}$ is the final reconstructed sequence. Since the token space is discrete and energy is derived from a sum of bounded logits, $E_{\theta}(X;Y)$ is lower-bounded by a finite minimum energy state. While stochastic sampling may introduce fluctuations, the diffusion process ensures progressive energy minimization, leading to an approximate low-energy state.
3.3.1 D-cMRF Guarantees
The above proof establishes that our method satisfies the following properties:
  • Progressive Energy Reduction: The energy function exhibits an overall decrease, leading to more stable sequence generation. This trend is supported by empirical results in Appendix D.
  • Stable Convergence: Since the energy function is lower-bounded and sequence length is finite, the generation process is expected to reach a structured, low-entropy state.
These properties explain the improved performance of Diffusion-EAGS compared to traditional diffusion models, as shown in §Section 6. Notably, the ablation study in Table 5 demonstrates that removing EAGS leads to a significant drop in performance, highlighting its importance in guiding stable generation.
4 Diffusion-EAGS
Our approach, Diffusion-EAGS, leverages two key components—Entropy-Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS)—rooted in the theory of Section 3. As shown in Figure 2, during training, ENS selectively masks tokens based on their certainty, while during generation, EAGS iteratively refines a fully masked sequence by updating high-uncertainty tokens. This stepwise refinement yields balanced improvements in text quality and diversity.
4.1 Inference Process: Entropy-Adaptive Gibbs Sampling
As discussed in Section 3.2, MLM can be interpreted as cMRF, which is used as $p_{\theta}$ in the sampling process of DDLM with $M_{t}$. In particular, $M_{t}$ is not only associated with energy updates but also serves as a solution to the MLM’s difficulty in selecting the next tokens to denoise, as shown in Appendix  C. Henceforth, we designate this strategy as Entropy-Adaptive Gibbs Sampling (EAGS).
In EAGS, $M_{t}$ is constructed by ranking tokens in descending order of entropy, thereby prioritizing the least informative parts of the sequence. EAGS facilitates the creation of more structured sequences by leveraging the syntactic context that has already been established. The process of determining the denoising schedule is shown in Appendix C.
Our approach for T step-generation process can be formalized as follows:
  1. 1.
    Entropy Calculation: Compute the entropy $H_{i}(X^{(t)})$ for each variable $x_{i}$.
  2. 2.
    Variable Selection: Obtain $M_{t}$ for sampling